Given: A(1,6), B(7,9), C(13,6), D(3,1).
Prove: ABCD is a trapazoid

Answer by shahid(44)

A trapezoid is a four sided geometrical figure having a pair of parallel and non-parallel sides.so in order to prove that givenfour points A(1,6),B(7,9),C(13,6)and D(3,1) form a trapezoid we have to prove that it has a pair of parallel sides and non-parallel sides.
now slope of AB m1=(y2-y1)/(x2-x1)
m1=9-6/7-1
m1=3/6
m1=1/2
similarly slope of CD m2=(1-6)/(3-13)
m2=-5/-10
m2=1/2
since m1=m2 so AB and CD are parallel sides similarly you can prove that sides AD and BC are non-parallel i-e their slopes are not equal.hope this will be helpfull